# Multi-Source EEG Emotion Recognition via Dynamic Contrastive Domain Adaptation

### v1

@@ III-C MS-DCDA Learning
where $N_{s}$ and $N_{t}$ denote the number of mini-batch features sampled respectively from features of individual source and the target, $N$ represent the total number of sources-target feature pairs and $k$ is the Guassian kernel function:
@@
Contrastive Sub-Domain Alignment. To improve the inter-class margin and improve the intra-class compactness of the alignment, we extend the contrastive domain discrepancy [36] to a sub-domain contrastive discrepancy (SCD) for the alignment of our multi-branch contrastive features. The SCD measures feature differences of pair-wise MBC features $\{f_{s}^{i},f_{t}^{i}\}$ in class level, which not only aligns inter-domain features in general but also aligns the intra-domain features with class-level accuracies:
@@
where $k(\cdot,\cdot)$ is the Gaussian kernal function (see Eq.3). The class masking function is given by:
@@
Classification and Feature Discrimination. Given the pair-wise multi-branch features $\{f_{s}^{i},f_{t}^{i}\}$, we train the classifier using cross entropy loss:
@@
where $P($·||||·)))) represents the probability distribution of the predicted label. Using the cross entropy loss alone tends to cause imbalanced prediction across different classifiers. To encourage a more consistent prediction of the target and accelerate the convergence rate of our multi-branch domain classifiers, a complementary classification loss $L_{DISC}$ is introduced:
$\mathcal{L}_{DISC}=\frac{1}{N^{2}}(\sum\limits_{j}^{N}\sum\limits_{i\neq j}^{N}E_{x\sim x_{t}}\mid\widehat{y}_{t}^{i}-\widehat{y}_{t}^{j}\mid)$ (11)
@@
$\arg\max\limits_{W}J(W)=\frac{Tr(W^{T}S_{b}W)}{Tr(W^{T}S_{W}W)}$ (12)
$\tau=\frac{L_{MMD}}{L_{MMD}+1-{J}(W)}$ (13)
$\mathcal{L}_{tot}=\mathcal{L}_{CE}+\alpha\big{(}(1-\tau)L_{DISC}+\tau\mathcal{L}_{MMD}\big{)}+\beta\mathcal{L}_{SCD}$ (14)
In addition to the dynamic coefficient $\tau$, another two dynamically changing coefficents include $\alpha$=($\frac{2}{1+e^{-10{\times}iter/epoch}}$)-1 and $\beta$=$\frac{\alpha}{10}$, where $iter$ is the index of iteration of training. The values of $\alpha$ and $\beta$ increase from 0 to 1 at different speeds as the number of iterations increases. The MS-DCDA algorithm is summarized in Algorithm 1.
@@
IV EXPERIMENTS
@@
TABLE I: CROSS-SUBJECT RESULTS OF A SINGLE ROUND OF LEAVE-ONE-OUT (LOO) VALIDATION
@@
DDC(2014) 72.78 88.57 62.61 81.59 77.02 99.56 81.76 72.33 87.57 82.82 73.90 82.76 81.14 94.52 78.93 81.19
DCORAL(2016) 81.65 76.78 79.32 75.63 79.88 81.70 73.10 79.55 74.93 82.47 80.26 76.84 85.21 75.37 75.63 78.56
DANN(2018) 71.77 86.53 61.67 81.26 78.29 99.77 77.34 72.01 93.02 77.28 73.40 83.50 81.85 87.10 83.03 80.52
DAN(2018) 84.74 85.48 80.73 76.16 87.63 87.86 79.02 82.74 81.56 87.86 80.00 75.93 80.14 81.29 81.47 82.17
MS-MDA(2021) 74.37 97.67 84.77 77.52 84.92 95.85 90.34 74.51 90.78 79.73 82.65 89.28 84.33 97.00 91.34 86.34
MS-ADA(2023) 87.74 91.22 100.00 77.34 95.49 80.08 82.15 79.41 86.01 98.68 83.44 86.01 89.45 79.52 89.51 87.07
Ours 76.58 98.70 75.99 79.11 90.51 100.00 83.53 88.21 100.00 88.80 95.58 99.29 92.55 94.87 99.00 90.85
@@
DDC(2014) 62.62 63.10 57.21 68.99 57.21 79.09 71.75 72.96 72.96 59.50 68.75 65.38 68.99 70.67 67.31 67.14
DCORAL(2016) 66.54 53.91 59.38 60.81 64.19 68.62 71.61 82.16 74.48 61.20 75.52 62.24 75.00 71.22 70.31 67.81
DANN(2018) 62.50 59.38 54.09 56.61 64.66 75.12 78.13 75.96 90.05 61.30 75.96 65.87 74.88 70.67 62.86 67.87
DAN(2018) 61.66 54.81 62.26 65.02 57.45 65.02 71.88 64.78 76.20 58.29 66.83 68.99 68.27 69.35 59.62 64.70
MS-MDA(2021) 74.52 52.28 67.43 74.04 72.48 78.37 91.47 91.11 85.70 71.63 75.00 66.47 91.47 77.52 72.12 76.11
MS-ADA(2023) 68.15 60.82 54.33 56.97 73.80 96.27 83.41 69.47 91.47 58.29 77.04 48.80 53.49 75.72 70.43 69.23
Ours 82.93 71.27 77.16 73.32 79.45 95.55 79.69 88.94 78.37 71.75 76.68 63.22 78.73 83.41 76.92 78.49
TABLE II: CROSS-SUBJECT RESULTS OF THREE LOO ROUNDS
@@
DDC(2014) 80.45 81.19±8.89 0.81 0.72
DCORAL(2016) 77.04 78.56±7.78 0.80 0.70
DANN(2018) 79.56 80.52±8.96 0.81 0.71
DAN(2018) 80.95 82.17±3.74 0.82 0.73
MS-MDA(2021) 85.74 86.34±7.43 0.86 0.79
MS-ADA(2023) 84.18 87.07±6.84 0.87 0.81
Ours 88.75 90.84±8.31 0.91 0.86
@@
DDC(2014) 65.22 67.14±6.03 0.62 0.54
DCORAL(2016) 65.72 67.81±7.29 0.65 0.57
DANN(2018) 63.78 67.87±8.14 0.67 0.58
DAN(2018) 57.64 64.70±5.62 0.59 0.51
MS-MDA(2021) 71.22 76.11±10.29 0.55 0.40
MS-ADA(2023) 65.95 69.23±13.62 0.67 0.57
Ours 74.34 78.49±7.36 0.79 0.71
IV-A Dataset
We evaluate baseline methods and our MS-DCDA model on the Shanghai Jiao Tong University (SJTU) Emotion EEG Dataset (SEED) [33][47] and the SJTU Emotion EEG Dataset IV (SEED-IV) [48]. Fifteen healthy participants (subjects), including seven males and eight females with an average age of 23 years, participated in data collection. The SEED dataset is collected by having participants watch 15 movie clips, which incite positive, neutral, and negative emotions. While subjects were watching movie clips, EEG data were obtained through a 62-channel ESI NeuroScan system. The collected data was downsampled to 200 Hz. The SEED signals are filtered with bandpass filter of 0-75Hz and segmented into non-overlapping intervals of 1 second each. The SEED-IV dataset is collected by having the same participants watch 24 movie clips, which incite neutral, sad, fear, and happy emotions. The SEED-IV signals are prefiltered by a 1–75 Hz bandpass filter and segmented into non-overlapping intervals of 4 seconds each. For both the SEED and SEED-IV dataset, signals are collected in three sessions for each suject and each segment.
Data Preprocessing. Each of the filtered EEG segments is considered as a data sample in training. DE features are extracted from each segment for five frequency bands: delta (1–4 Hz), theta (4–8 Hz), alpha (8–14 Hz), beta (14–31 Hz), and gamma (31–50 Hz), resulting in a feature dimension of 62 channels × 5 frequency bands. A total of 3394 EEG samples are collected for each participant per session for the SEED set. The SEED-IV set consists of 851, 832, 822 samples per participant for each of the three sessions respectively. Labels are generated for both datasets, and pre-processed EEG data of each participant are normalized electrode-wise [49] based on individual mean and standard deviation values.
TABLE III: CROSS-SUBJECT RESULTS OF THREE LOO ROUNDS
COMPARISON WITH MORE METHODS (ACC-best)
@@ IV-A Dataset
BiDANN(2018) [50] 84.14±6.87 65.59±10.39
RGNN(2020) [51] 85.30±6.72 73.84±8.02
TANN(2021) [52] 84.41±8.75 68.00±8.35
PPDA(2021) [53] 86.70±7.10 -
GMSS(2022) [54] 86.52±6.22 73.48±7.41
HVF2N-DBR(2022) [55] 89.33±10.13 73.60±2.91
MWACN(2022) [15] 89.30±9.18 74.60±10.77
PCDG(2023) [56] 87.30±2.10 73.60±5.10
MFA-LR(2023) [44] 89.11±7.72 74.99±12.1
Ours 90.84±8.31 78.49±7.36
IV-B Experimental Setup
To evaluate the performance of the proposed MS-DCDA network, we performed three round leave-one-out (LOO) cross-validations for cross-subject and cross-session experiments respectively. In the Cross-Subject experiment, a total number of 15 subjects is evaluated, where EEG of one subject is selected as the target and EEG of the remaining 14 subjects are considered as sources. EEG signals of subjects measured in one session are selected as target, and EEG signals of subjects measured from the rest two sessions are considered as source.
IV-C Training
The proposed MS-DCDA model is trained with Adam optimizer for 50 epochs. We use a batch size of 32 and 16 respectively for the SEED and SEED-IV dataset. The learning rate remains at 5 × $10^{-3}$ for both. The dynamic coefficients are given by $\alpha$=($\frac{2}{1+e^{-10{\times}iter/epoch}}$)-1 and $\beta$=$\frac{\alpha}{10}$ respectively, where $iter$ is the index of training iterations. All the training tasks are conducted on a NVIDIA GeForce RTX 3060 GPU. The implementation is based on Python 3.9.18, PyTorch 2.0.0, and Torchvision 0.15.0.
TABLE IV: CROSS-SESSION RESULTS OF A SINGLE ROUND OF LEAVE-ONE-OUT (LOO) VALIDATION
@@ IV-C Training
DDC(2014) 93.64 94.99 91.04 96.67 93.70 98.85 78.82 90.19 83.32 81.67 76.78 92.55 90.51 97.17 87.13 89.80
DCORAL(2016) 82.21 93.03 90.90 87.26 84.19 97.66 83.05 89.63 80.32 85.58 77.82 90.02 90.90 97.51 79.39 87.30
DANN(2018) 90.36 96.14 91.42 96.02 95.28 97.97 79.13 90.74 81.90 84.94 78.01 89.33 90.45 99.00 80.84 89.43
DAN(2018) 91.42 100.00 90.92 92.39 98.70 100.00 82.25 95.25 85.97 82.16 81.10 100.00 92.13 99.73 89.00 92.07
MS-MDA(2021) 94.43 90.78 91.34 100.00 100.00 100.00 84.12 97.17 92.78 92.13 84.03 91.13 96.61 97.61 91.28 93.61
MS-ADA(2023) 84.55 93.12 98.44 92.10 97.73 100.00 83.87 100.00 98.11 91.27 79.36 85.55 95.05 100.00 100.00 93.28
Ours 88.51 90.75 100.00 90.95 100.00 93.96 100.00 99.73 84.68 95.82 99.53 95.76 99.03 100.00 98.59 95.82
@@
DDC(2014) 74.75 59.88 66.13 61.88 92.50 90.25 81.50 81.75 57.75 84.25 86.75 91.88 91.50 83.50 66.00 78.02
DCORAL(2016) 74.74 66.41 64.71 73.44 95.44 82.81 71.22 82.29 54.04 85.68 76.69 86.33 89.58 79.82 56.90 76.01
DANN(2018) 73.32 51.32 68.27 76.20 87.02 94.23 66.47 80.89 83.17 78.61 69.47 79.57 90.02 73.68 57.57 75.32
DAN(2018) 70.13 59.63 45.13 73.50 92.25 86.75 78.50 91.00 56.88 82.88 74.63 93.00 92.38 80.13 64.63 76.09
MS-MDA(2021) 75.36 51.80 73.92 85.46 78.00 93.15 71.51 87.02 85.46 78.00 81.97 72.36 85.82 86.78 72.36 78.60
MS-ADA(2023) 64.13 56.63 61.00 60.13 98.00 93.38 76.75 81.88 60.88 64.63 78.88 86.13 100.00 82.50 77.63 76.17
Ours 81.37 73.16 69.24 66.79 97.92 92.65 72.43 94.98 68.01 79.53 84.68 90.56 95.22 94.85 72.30 82.25
TABLE V: CROSS-SESSION RESULTS OF THREE LOO ROUNDS
@@
DDC(2014) 89.15 89.80±6.61 0.88 0.81
DCORAL(2016) 85.57 87.30±6.00 0.87 0.81
DANN(2018) 88.14 89.43±6.74 0.88 0.81
DAN(2018) 90.60 92.07±6.66 0.94 0.90
MS-MDA(2021) 91.80 93.61±4.95 0.94 0.91
MS-ADA(2023) 92.02 93.28±6.72 0.93 0.89
Ours 93.92 95.82±4.81 0.95 0.93
@@
DDC(2014) 75.30 78.02±12.11 0.77 0.70
DCORAL(2016) 72.86 76.01±11.44 0.74 0.67
DANN(2018) 72.97 75.32±11.23 0.74 0.67
DAN(2018) 72.79 76.09±14.12 0.74 0.68
MS-MDA(2021) 75.78 78.60±9.61 0.73 0.66
MS-ADA(2023) 72.22 76.17±13.91 0.77 0.69
Ours 77.53 82.25±11.00 0.83 0.77
V RESULTS
Three rounds of LOO cross-validations (see Section IV-B) are conducted for cross-subject and cross-session experiments respectively. The same validations are repeated for the SEED and SEED-IV datasets respectively. For each dataset and experiment, we evaluated the mean recognition accuracy (ACC-mean) and the best recognition accuracy (ACC-best) across the three LOO validation rounds for the target domain. The best recognition accuracies are given in the form of mean $\pm$ standard deviation for the three rounds of LOO.
We compare the proposed MS-DCDA with several advanced methods for both cross-subject and cross-session experiments. They include four single-source domain adaptation methods: deep domain confusion (DDC) [57], DAN [35], DANN [36], deep correlation alignment (DCORAL) [58]; as well as two multi-source domain adaptation models: multisource marginal distribution adaptation (MS-MDA) [49], and multi-source associate domain adaptation (MS-ADA) [43]. The DDC method minimizes the domain distribution discrepancies between the source and the target using MMD. The DCORAL model aligns the second-order statistics of domain distributions using linear transformations. The DANN aligns domain feature distributions through backpropagation. The DAN method aligns high-level domain features through the use of deep neural networks and multi-kernel MMD. In the MS-MDA method, multi-branch networks and MMD are combined for the multi-source alignment. The MS-ADA aligns the edge distributions and through MMD and reinforcement learning respectively. In addition, we compare our MS-DCDA against several more models for the cross-subject experiment. They include BiDANN [50], RGNN [51], TANN [52], PPDA [53], GMSS [54], HVF2N-DBR [55], MWACN [15], PCDG [56], and MFA-LR [44].
V-A Cross-Subject Results

### v2

@@ 3.3 MS-DCDA Learning
where $N_{s}$ and $N_{t}$ denote the number of mini-batch features sampled respectively from features of individual source and the target, $N$ represent the total number of sources-target feature pairs and $k$ is the Gaussian kernel function:
@@
Contrastive Sub-Domain Alignment. To improve the inter-class margin and improve the intra-class compactness of the alignment, we extend the contrastive domain discrepancy [42] to a sub-domain contrastive discrepancy (SCD) for the alignment of our multi-branch contrastive features. The SCD measures feature differences of pair-wise MBC features $\{f_{s,D}^{i},f_{t,D}^{i}\}$ in class level, which not only aligns inter-domain features in general but also aligns the intra-domain features with class-level accuracies:
@@
where $k(\cdot,\cdot)$ is the Gaussian kernel function (see Eq.3). The class masking function is given by:
@@
Classification and Feature Discrimination. Given the pair-wise multi-branch features $\{f_{s,D}^{i},f_{t,D}^{i}\}$, we train the classifier using cross entropy loss:
@@
where $P($·||||·)))) represents the probability distribution of the predicted label. Using the cross entropy loss alone typically leads to imbalanced prediction across different classifiers. To encourage consistent prediction of the target and accelerate the convergence rate of our multi-branch domain classifiers, a complementary classification loss $L_{DISC}$ is introduced:
$\mathcal{L}_{DISC}=\frac{1}{N}\bigg{(}\sum\limits_{i\neq j}^{N}E_{x\sim x_{t}}\mid\widehat{y}_{t}^{i}-\widehat{y}_{t}^{j}\mid\bigg{)}$ (11)
@@
$\arg\max\limits_{W}J(W)=\frac{Tr(W^{T}S_{b}W)}{Tr(W^{T}S_{W}W)}$ (12)
$\tau=\frac{L_{MMD}}{L_{MMD}+1-{J}(W)}$ (13)
$\mathcal{L}_{tot}=\mathcal{L}_{CE}+\alpha\cdot\big{(}(1-\tau)L_{DISC}+\tau\mathcal{L}_{MMD}\big{)}+\beta\cdot\mathcal{L}_{SCD}$ (14)
In addition to the dynamic coefficient $\tau$, another two dynamically changing coefficients include $\alpha$=($\frac{2}{1+e^{-10{\times}iter/epoch}}$)-1 and $\beta$=$\frac{\alpha}{10}$, where $iter$ is the index of iteration of training. The values of $\alpha$ and $\beta$ increase from 0 to 1 at different speeds as the number of iterations increases. The MS-DCDA algorithm is summarized in Algorithm 1.
4 EXPERIMENTS
@@
4.1 Dataset
We evaluate baseline methods and our MS-DCDA model on the Shanghai Jiao Tong University (SJTU) Emotion EEG Dataset (SEED) [34][54] and the SJTU Emotion EEG Dataset IV (SEED-IV) [55]. Fifteen healthy participants (subjects), including seven males and eight females with an average age of 23 years, participated in data collection. The SEED dataset is collected by having participants watch 15 movie clips, which incite positive, neutral, and negative emotions. While subjects were watching movie clips, EEG data were obtained through a 62-channel ESI NeuroScan system. The collected data was downsampled to 200 Hz. The SEED signals are filtered with bandpass filter of 0-75Hz and segmented into non-overlapping intervals of 1 second each. The SEED-IV dataset is collected by having the same participants watch 24 movie clips, which incite neutral, sad, fear, and happy emotions. The SEED-IV signals are prefiltered by a 1-75 Hz bandpass filter and segmented into non-overlapping intervals of 4 seconds each. For both the SEED and SEED-IV dataset, signals are collected in three sessions for each subject and each segment.
Data Preprocessing. Each of the filtered EEG segments is considered as a data sample in training. DE features are extracted from each segment for five frequency bands: delta (1-4 Hz), theta (4-8 Hz), alpha (8-14 Hz), beta (14-31 Hz), and gamma (31-50 Hz), resulting in a feature dimension of 62 channels $\times$ 5 frequency bands. A total of 3394 EEG samples are collected for each participant per session for the SEED set. The SEED-IV set consists of 851, 832, 822 samples per participant for each of the three sessions respectively. Labels are generated for both datasets, and pre-processed EEG data of each participant are normalized electrode-wise [56] based on individual mean and standard deviation values.
4.2 Experimental Setup
To evaluate the performance of the proposed MS-DCDA network, we performed three round leave-one-out (LOO) cross validations for cross-subject and cross-session experiments respectively. In the Cross-Subject experiment, a total number of 15 subjects is evaluated, where EEG of one subject is selected as the target and EEG of the remaining 14 subjects are considered as sources. EEG signals of subjects measured in one session are selected as target, and EEG signals of subjects measured from the rest two sessions are considered as source.
4.3 Training
The proposed MS-DCDA model is trained with Adam optimizer for 50 epochs. We use a batch size of 32 and 16 respectively for the SEED and SEED-IV dataset. The learning rate remains at 5 $\times$ $10^{-3}$ for both. The dynamic coefficients are given by $\alpha$=($\frac{2}{1+e^{-10{\times}iter/epoch}}$)-1 and $\beta$=$\frac{\alpha}{10}$ respectively, where $iter$ is the index of training iterations. All the training tasks are conducted on a NVIDIA GeForce RTX 3060 GPU. The implementation is based on Python 3.9.18, PyTorch 2.0.0, and Torchvision 0.15.0.
Table 1: Cross-subject Results of a Single Round of Leave-one-out (LOO) Validation
@@ 4.3 Training
DDC [57] 72.78 88.57 62.61 81.59 77.02 99.56 81.76 72.33 87.57 82.82 73.90 82.76 81.14 94.52 78.93 81.19
DCORAL [58] 81.65 76.78 79.32 75.63 79.88 81.70 73.10 79.55 74.93 82.47 80.26 76.84 85.21 75.37 75.63 78.56
DANN [42] 71.77 86.53 61.67 81.26 78.29 99.77 77.34 72.01 93.02 77.28 73.40 83.50 81.85 87.10 83.03 80.52
DAN [41] 84.74 85.48 80.73 76.16 87.63 87.86 79.02 82.74 81.56 87.86 80.00 75.93 80.14 81.29 81.47 82.17
MS-MDA [56] 74.37 97.67 84.77 77.52 84.92 95.85 90.34 74.51 90.78 79.73 82.65 89.28 84.33 97.00 91.34 86.34
MS-ADA [50] 87.74 91.22 100.00 77.34 95.49 80.08 82.15 79.41 86.01 98.68 83.44 86.01 89.45 79.52 89.51 87.07
MS-DCDA (ours) 76.58 98.70 75.99 79.11 90.51 100.00 83.53 88.21 100.00 88.80 95.58 99.29 92.55 94.87 99.00 90.85
@@
DDC [57] 62.62 63.10 57.21 68.99 57.21 79.09 71.75 72.96 72.96 59.50 68.75 65.38 68.99 70.67 67.31 67.14
DCORAL [58] 66.54 53.91 59.38 60.81 64.19 68.62 71.61 82.16 74.48 61.20 75.52 62.24 75.00 71.22 70.31 67.81
DANN [42] 62.50 59.38 54.09 56.61 64.66 75.12 78.13 75.96 90.05 61.30 75.96 65.87 74.88 70.67 62.86 67.87
DAN [41] 61.66 54.81 62.26 65.02 57.45 65.02 71.88 64.78 76.20 58.29 66.83 68.99 68.27 69.35 59.62 64.70
MS-MDA [56] 74.52 52.28 67.43 74.04 72.48 78.37 91.47 91.11 85.70 71.63 75.00 66.47 91.47 77.52 72.12 76.11
MS-ADA [50] 68.15 60.82 54.33 56.97 73.80 96.27 83.41 69.47 91.47 58.29 77.04 48.80 53.49 75.72 70.43 69.23
MS-DCDA (ours) 82.93 71.27 77.16 73.32 79.45 95.55 79.69 88.94 78.37 71.75 76.68 63.22 78.73 83.41 76.92 78.49
Table 2: Cross-Subject Results of Three LOO Rounds
@@
DDC [57] 80.45 81.19$\pm$8.89 0.81 0.72
DCORAL [58] 77.04 78.56$\pm$7.78 0.80 0.70
DANN [42] 79.56 80.52$\pm$8.96 0.81 0.71
DAN [41] 80.95 82.17$\pm$3.74 0.82 0.73
MS-MDA [56] 85.74 86.34$\pm$7.43 0.86 0.79
MS-ADA [50] 84.18 87.07$\pm$6.84 0.87 0.81
MS-DCDA (ours) 88.75 90.84$\pm$8.31 0.91 0.86
@@
DDC [57] 65.22 67.14$\pm$6.03 0.62 0.54
DCORAL [58] 65.72 67.81$\pm$7.29 0.65 0.57
DANN [42] 63.78 67.87$\pm$8.14 0.67 0.58
DAN [41] 57.64 64.70$\pm$5.62 0.59 0.51
MS-MDA [56] 71.22 76.11$\pm$10.29 0.55 0.40
MS-ADA [50] 65.95 69.23$\pm$13.62 0.67 0.57
MS-DCDA (ours) 74.34 78.49$\pm$7.36 0.79 0.71
Table 3: Cross-Subject Results of Three LOO Rounds (ACC-best)
@@
       BiDANN [59] 84.14$\pm$6.87 65.59$\pm$10.39
       RGNN [60] 85.30$\pm$6.72 73.84$\pm$8.02
       TANN [61] 84.41$\pm$8.75 68.00$\pm$8.35
       PPDA [62] 86.70$\pm$7.10 -
       GMSS [63] 86.52$\pm$6.22 73.48$\pm$7.41
       HVF2N-DBR [64] 89.33$\pm$10.13 73.60$\pm$2.91
       MWACN [15] 89.30$\pm$9.18 74.60$\pm$10.77
       PCDG [65] 87.30$\pm$2.10 73.60$\pm$5.10
       MFA-LR [39] 89.11$\pm$7.72 74.99$\pm$12.1
       MS-DCDA (ours) 90.84$\pm$8.31 78.49$\pm$7.36
5 RESULTS
Three rounds of LOO cross-validations (see Section 4.2) are conducted for cross-subject and cross-session experiments respectively. The same validations are repeated for the SEED and SEED-IV datasets respectively. For each dataset and experiment, we evaluated the mean recognition accuracy (ACC-mean) and the best recognition accuracy (ACC-best) across the three LOO validation rounds for the target domain. The best recognition accuracies are given in the form of mean $\pm$ standard deviation for the three rounds of LOO.
We compare the proposed MS-DCDA with several advanced methods for both cross-subject and cross-session experiments. They include four single-source domain adaptation methods: deep domain confusion (DDC) [57], DAN [41], DANN [42], deep correlation alignment (DCORAL) [58]; Two multi-source domain adaptation models: multisource marginal distribution adaptation (MS-MDA) [56], and multi-source associate domain adaptation (MS-ADA) [50]. The DDC method minimizes the domain distribution discrepancies between the source and the target using MMD. The DCORAL model aligns the second-order statistics of domain distributions using linear transformations. The DANN aligns domain feature distributions through backpropagation. The DAN method aligns high-level domain features through the use of deep neural networks and multi-kernel MMD. In the MS-MDA method, multi-branch networks and MMD are combined for the multi-source alignment. The MS-ADA aligns the edge distributions and through MMD and reinforcement learning respectively. In addition, we compare our MS-DCDA against several more models for the cross-subject experiment. They include BiDANN [59], RGNN [60], TANN [61], PPDA [62], GMSS [63], HVF2N-DBR [64], MWACN [15], PCDG [65], and MFA-LR [39].
Table 4: Cross-session Results of a Single Round of Leave-one-out (LOO) Validation
@@ 5 RESULTS
DDC [57] 93.64 94.99 91.04 96.67 93.70 98.85 78.82 90.19 83.32 81.67 76.78 92.55 90.51 97.17 87.13 89.80
DCORAL [58] 82.21 93.03 90.90 87.26 84.19 97.66 83.05 89.63 80.32 85.58 77.82 90.02 90.90 97.51 79.39 87.30
DANN [42] 90.36 96.14 91.42 96.02 95.28 97.97 79.13 90.74 81.90 84.94 78.01 89.33 90.45 99.00 80.84 89.43
DAN [41] 91.42 100.00 90.92 92.39 98.70 100.00 82.25 95.25 85.97 82.16 81.10 100.00 92.13 99.73 89.00 92.07
MS-MDA [56] 94.43 90.78 91.34 100.00 100.00 100.00 84.12 97.17 92.78 92.13 84.03 91.13 96.61 97.61 91.28 93.61
MS-ADA [50] 84.55 93.12 98.44 92.10 97.73 100.00 83.87 100.00 98.11 91.27 79.36 85.55 95.05 100.00 100.00 93.28
MS-DCDA (ours) 88.51 90.75 100.00 90.95 100.00 93.96 100.00 99.73 84.68 95.82 99.53 95.76 99.03 100.00 98.59 95.82
@@
DDC [57] 74.75 59.88 66.13 61.88 92.50 90.25 81.50 81.75 57.75 84.25 86.75 91.88 91.50 83.50 66.00 78.02
DCORAL [58] 74.74 66.41 64.71 73.44 95.44 82.81 71.22 82.29 54.04 85.68 76.69 86.33 89.58 79.82 56.90 76.01
DANN [42] 73.32 51.32 68.27 76.20 87.02 94.23 66.47 80.89 83.17 78.61 69.47 79.57 90.02 73.68 57.57 75.32
DAN [41] 70.13 59.63 45.13 73.50 92.25 86.75 78.50 91.00 56.88 82.88 74.63 93.00 92.38 80.13 64.63 76.09
MS-MDA [56] 75.36 51.80 73.92 85.46 78.00 93.15 71.51 87.02 85.46 78.00 81.97 72.36 85.82 86.78 72.36 78.60
MS-ADA [50] 64.13 56.63 61.00 60.13 98.00 93.38 76.75 81.88 60.88 64.63 78.88 86.13 100.00 82.50 77.63 76.17
MS-DCDA (ours) 81.37 73.16 69.24 66.79 97.92 92.65 72.43 94.98 68.01 79.53 84.68 90.56 95.22 94.85 72.30 82.25
Table 5: Cross-Session Results of Three Leave-one-out (LOO) Rounds
@@
DDC [57] 89.15 89.80$\pm$6.61 0.88 0.81
DCORAL [58] 85.57 87.30$\pm$6.00 0.87 0.81
DANN [42] 88.14 89.43$\pm$6.74 0.88 0.81
DAN [41] 90.60 92.07$\pm$6.66 0.94 0.90
MS-MDA [56] 91.80 93.61$\pm$4.95 0.94 0.91
MS-ADA [50] 92.02 93.28$\pm$6.72 0.93 0.89
MS-DCDA(ours) 93.92 95.82$\pm$4.81 0.95 0.93
@@
DDC [57] 75.30 78.02$\pm$12.11 0.77 0.70
DCORAL [58] 72.86 76.01$\pm$11.44 0.74 0.67
DANN [42] 72.97 75.32$\pm$11.23 0.74 0.67
DAN [41] 72.79 76.09$\pm$14.12 0.74 0.68
MS-MDA [56] 75.78 78.60$\pm$9.61 0.73 0.66
MS-ADA [50] 72.22 76.17$\pm$13.91 0.77 0.69
MS-DCDA (ours) 77.53 82.25$\pm$11.00 0.83 0.77
5.1 Cross-Subject Results